# Time-Dependent Black Hole Lensing and the Unified Weak-to-Strong Deflection Framework

### v1

@@ Appendix B Derivation of the master Born kernel
Our goal is to derive the line-of-sight expression that underlies Eqs. (19)-(24) for the screen-plane deflection produced by a first-order metric perturbation $h_{\mu\nu}$ on the Schwarzschild background of Eq. (1). We keep the perturbative split of Eq. (3), use the observer tetrad of Eq. (9), and map momenta to screen angles via Eq. (13). Throughout, we follow a Born, or single-scattering, treatment: the null ray follows the unperturbed geodesic $x_{0}^{\mu}(\lambda;b)$ while its momentum $p^{\mu}$ is corrected at $\mathcal{O}(\varepsilon)$.
@@
$\dot{x}_{0}^{\mu}=\frac{\partial H_{0}}{\partial p_{0\mu}}=g^{(0)\mu\nu}p_{0\nu}\equiv p_{0}^{\mu},\qquad\dot{p}_{0\mu}=-\frac{\partial H_{0}}{\partial x_{0}^{\mu}}=0,$ (79)
@@
$\dot{\delta p}_{\mu}=-\frac{\partial}{\partial x^{\mu}}\,\left(\frac{1}{2}\,h^{\alpha\beta}p_{0\alpha}p_{0\beta}\right)=-\frac{1}{2}\,\partial_{\mu}h_{\alpha\beta}\,p_{0}^{\alpha}p_{0}^{\beta},$ (80)
which integrates to Eq. (18) once evaluated between source $\lambda_{s}$ and observer $\lambda_{o}$:
$\delta p_{\mu}(\lambda_{o};b)=-\frac{1}{2}\int_{\lambda_{s}}^{\lambda_{o}}\,d\lambda\,\partial_{\mu}h_{\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}(\lambda)\,p_{0}^{\beta}(\lambda).$ (81)
Eq. (81) is the covariant Born response on a fixed background path.
Two comments are useful. First, the “force” driving $\delta p_{\mu}$ can equivalently be written in terms of the perturbed Christoffel symbols, since
@@
$\delta\Gamma_{\mu\alpha\beta}=\frac{1}{2}\,\left(\nabla_{\alpha}h_{\mu\beta}+\nabla_{\beta}h_{\mu\alpha}-\nabla_{\mu}h_{\alpha\beta}\right),$ (82)
so that integrating by parts turns $B.4$ into a $\delta\Gamma$-driven integral plus boundary terms. Second, the contraction $h_{\alpha\beta}p_{0}^{\alpha}p_{0}^{\beta}$ used below is given in (72)-(75) in terms of the Zerilli master field $\Psi_{Z}$.
@@
$p^{\hat{a}}=e^{\hat{a}}{}_{\mu}p^{\mu}=p_{0}^{\hat{a}}+\varepsilon\,\delta p^{\hat{a}},\qquad\delta p^{\hat{a}}=e^{\hat{a}}{}_{\mu}\,\delta p^{\mu},$ (83)
@@
$\alpha_{A}$ $\equiv\frac{A}{r_{o}}=\frac{p^{\widehat{A}}}{p^{\hat{0}}},\quad A\in{X,Y}.$ (84)
Varying Eq. (84) and keeping terms linear in $\varepsilon$ gives
$\delta\alpha_{A}=\frac{\delta p^{\widehat{A}}}{p_{0}^{\hat{0}}}-\frac{p_{0}^{\widehat{A}}}{(p_{0}^{\hat{0}})^{2}}\,\delta p^{\hat{0}}=\frac{1}{p_{0}^{\hat{0}}}\,\Pi_{A}{}^{\ \hat{b}}\,\delta p_{\hat{b}}$, (85)
@@
$\Pi_{A}{}^{\ \hat{b}}=\delta_{A}{}^{\ \hat{b}}-\alpha_{0,A}\,\delta^{\hat{b}}{}_{\hat{0}}\,,\qquad\alpha_{0,A}=\frac{p_{0}^{\widehat{A}}}{p_{0}^{\hat{0}}}$, (86)
and hatted indices are raised/lowered with $\eta_{\hat{a}\hat{b}}$. Using $\delta p_{\hat{b}}=e_{\hat{b}}{}^{\mu}\,\delta p_{\mu}$ and substituting Eq. (81) then yields
@@
$\times\partial_{\mu}h_{\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}(\lambda)p_{0}^{\beta}(\lambda),$ (87)
which is Eq. (20) with the explicit kernel
@@
$\times\partial_{\mu}h_{\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}p_{0}^{\beta}.$ (88)
Eq. (88) is identical to Eq. (21) once we identify the tetrad time leg $e^{\nu}{}_{\hat{0}}$ and note that $\Pi_{A}{}^{\ \mu}=\Pi_{A}{}^{\ \hat{b}}e_{\hat{b}}{}^{\ \mu}$.
Equivalent kernel forms and boundary terms: Using the identity Eq. (82) and integrating by parts, Eq. (87) becomes
$\delta\alpha_{A}=\int_{\lambda_{s}}^{\lambda_{o}}\,d\lambda\,\mathcal{K}^{(\Gamma)}_{A}+\mathcal{B}_{A}\Big|_{\lambda_{s}}^{\lambda_{o}},$ (89)
@@
$\mathcal{K}^{(\Gamma)}_{A}=-\frac{1}{p_{0}^{\hat{0}}(\lambda_{o})}\,\Pi_{A}{}^{\ \hat{b}}\,e_{\hat{b}}{}^{\ \mu}(x_{o})\,\delta\Gamma_{\mu\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}p_{0}^{\beta},$
$\mathcal{B}_{A}=\frac{1}{2\,p_{0}^{\hat{0}}(\lambda_{o})}\,\Pi_{A}{}^{\ \hat{b}}\,e_{\hat{b}}{}^{\ \mu}(x_{o})\,h_{\mu\nu}\left(x_{0}(\lambda)\right)\,p_{0}^{\nu}.$ (90)
@@
The boundary term $\mathcal{B}_{A}$ vanishes under our asymptotic conditions: at the observer, the screen projector eliminates any $\hat{0}$ contamination and $h_{\mu\nu}(x_{o})=\mathcal{O}(r_{o}^{-1})$ [Eq. (71)]; at the source, we assume compact support or sufficiently rapid falloff. Hence, all kernel representatives obtained by redistributing derivatives are equivalent at $\mathcal{O}(\varepsilon)$.
@@
A further useful form follows by changing variables from affine parameter $\lambda$ to the tortoise coordinate $r_{*}$ or to the delay $\tau$ of Eq. (24). Using Eq. (75),
@@
$\delta\alpha_{A}=\int_{\tau=0}^{\infty}\,d\tau\,\widetilde{\mathcal{K}}_{A}\left[r(\tau),\theta(\tau)\right]\,e^{i\omega\tau}\times e^{-i\omega t_{o}}+\text{c.c.},$ (91)
@@
which is the factorized form quoted in Eq. (24) once the QNM time dependence in Eq. (23) is inserted.
@@
Gauge behavior at $\mathcal{O}(\varepsilon)$: Under a linearized diffeomorphism generated by $\xi^{\mu}$, Eq. (15) gives $\delta h_{\alpha\beta}=-\nabla_{\alpha}\xi_{\beta}-\nabla_{\beta}\xi_{\alpha}$. Substituting this into Eq. (87) and integrating by parts yields
@@
$\delta_{\xi}\alpha_{A}=\frac{1}{p_{0}^{\hat{0}}(\lambda_{o})}\,\Pi_{A}{}^{\ \hat{b}}\,e_{\hat{b}}{}^{\ \mu}(x_{o})\,\xi_{\mu}\Big|_{\lambda_{s}}^{\lambda_{o}}+\mathcal{O}\,\left(\frac{\varepsilon}{r_{o}}\right),$ (92)
@@
where we used the background geodesic equation and $\dot{p}_{0\mu}=0$. With our asymptotic conditions $\xi^{\mu}=\mathcal{O}(1/r)$, Eq. (92) reproduces the gauge-independence statement in Eq (16): $\delta_{\xi}\alpha_{A}=0+\mathcal{O}(\varepsilon/r_{o})$. Thus, the kernel Eq. (88) produces gauge-safe screen observables at first order, provided the observer tetrad is fixed by Eq. (10).
@@
Far-zone reduction and the (1/b) law: Insert the large-$r$ asymptotics $A.12$ into $B.11$, parametrize the background ray by a straight line with longitudinal coordinate $z$ as in Eq. (26), and note that $p_{0}^{\hat{0}}(\lambda_{o})=E$ to leading order. One finds
@@
$\delta\alpha_{A}(b,t_{o})=\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{c_{A}(\omega)}{b}\,\mathcal{F}_{A}(\omega b)\right\}+\mathcal{O}\,\left(\frac{M}{b^{2}}\right),$ (93)
@@
with $c_{A}$ and $\mathcal{F}_{A}$ defined in Eqs. (27)-(28). This reproduces the weak-field Born expansion used in Sec. III.2 and provides the link between the kernel normalization and the centroid wobble of Eq. (32).
@@
Near-photon-sphere reduction and the logarithmic law: In the near-critical regime $b\to b_{c}^{+}$, reparametrize the background ray by the azimuth $\varphi$ accumulated during the whirl [Eq. (36)]. Use the near-photon-sphere reduction $A.13$ inside $B.11$ and integrate over the long dwell time $\Delta\varphi\sim-\ln\Delta$, with $\Delta=b/b_{c}-1$. The integral yields
@@
$\delta\alpha_{A}(b,t_{o})=\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\left[\mathcal{C}_{A}(\omega)\,\ln\Delta+\mathcal{D}_{A}(\omega)\right]\right\}+\ldots,$ (94)
@@
which is Eq. (37). When combined with the background divergence in Eq. (34) and reorganized as in Eqs. (38)-(41), Eq. (94) fixes the time-dependent SDL coefficients $a_{1},b_{1},\beta_{1}$ from the same kernel.
@@
Starting from Hamilton’s equations, we derived the Born response (81) for the photon momentum, projected it to the observer’s screen to obtain the master kernel (88), and exhibited equivalent $\delta\Gamma$ and delay-space forms, Eqs. (90)-(91). The kernel is gauge-safe at $\mathcal{O}(\varepsilon)$ by Eq. (92). Its far-zone and near-photon-sphere reductions, Eqs. $B.16$ and $B.17$, reproduce the weak-field $1/b$ behavior Sec. III.2 and the logarithmic SDL structure Sec. IV, respectively. In this sense, all imaging diagnostics discussed in Sec. V are different asymptotic faces of a single line-of-sight kernel driven by the Zerilli master field.
@@ Appendix C Near-critical expansion and SDL coefficients
In this appendix, we provide a self-contained derivation of the near-critical expansion underlying Eqs. (34) and (38)-(41), and we express the time-dependent strong-deflection-limit $SDL$ response coefficients $a_{1}(\omega),b_{1}(\omega),\beta_{1}(\omega)$} in terms of quantities localized at the photon sphere and of moments of the Born kernel introduced in Appendix B. We keep all conventions of Secs. II-IV, in particular the Schwarzschild background (Eq. (1)), the perturbative split (Eq. (3)), and the master Born representation (Eqs. (19)-(24)).
@@
$\mathcal{F}(r)\equiv r\,A^{\prime}(r)-2A(r)=0,\qquad r=r_{\mathrm{ph}},$ (96)
and the critical impact parameter is $b_{c}=r_{\mathrm{ph}}/\sqrt{A(r_{\mathrm{ph}})}$. For Schwarzschild, A=f=12M/rA=f=1-2M/r gives $r_{\mathrm{ph}}=3M$ and $b_{c}=3\sqrt{3}M$ as stated in Eq. (33). Expanding the background bending integral $16$ around $b\to b_{c}^{+}$ yields Eq. (34),
$\alpha_{0}(b)=-\bar{a}\,\ln\,\left(\frac{b}{b_{c}}-1\right)+\bar{b}+\mathcal{O}\,\left((b-b_{c})\ln|b-b_{c}|\right),,$ (97)
with $\bar{b}$ is a constant depending on the chosen normalization, and $\bar{a}=1$ for the Schwarzschild case.
Linear response of the critical scale $b_{c}(t_{o})$: We now compute the $\mathcal{O}(\varepsilon)$ modulation of $b_{c}$ induced by the even-parity $(\ell,m)=(2,0)$ perturbation described in Sec. II.1. We adopt an “instantaneous snapshot” characterization at the detection time $t_{o}$: the time dependence appears as a harmonic factor $e^{-i\omega t_{o}}$ multiplying complex amplitudes evaluated at $r_{\mathrm{ph}}$. This adiabatic reading is equivalent, to first order, to the near-photon-sphere reduction of the Born kernel and provides a compact route to $\beta_{1}(\omega)$.
Let $A(r)\to A(r)+\varepsilon\,a(r,t)$, where $A=f$ for the background and $a$ is the even-parity perturbation of $g_{tt}$ (we keep $B$ unmodified since, to first order in $\varepsilon$, $b_{c}$ is controlled by $A$ alone through Eq. (95)). The photon-sphere condition (96) becomes $\mathcal{F}(r_{\mathrm{ph}}+\varepsilon\,\delta r)=0$ with
$\delta r\,\mathcal{F}^{\prime}(r_{\mathrm{ph}})+\left[r\,a^{\prime}(r)-2a(r)\right]_{r=r_{\mathrm{ph}}}=0,$
$\mathcal{F}^{\prime}(r_{\mathrm{ph}})=r_{\mathrm{ph}}A^{\prime\prime}(r_{\mathrm{ph}})-A^{\prime}(r_{\mathrm{ph}}).$ (98)
Solving Eq. (98) gives
@@
$=-\frac{3M}{2}\,\left[2\,a(r_{\mathrm{ph}})-r_{\mathrm{ph}}\,a^{\prime}(r_{\mathrm{ph}})\right].$ (99)
@@
$\frac{\delta b_{c}}{b_{c}}=\frac{\delta r}{r_{\mathrm{ph}}}-\frac{1}{2}\left[\frac{a(r_{\mathrm{ph}})}{A(r_{\mathrm{ph}})}+\frac{A^{\prime}(r_{\mathrm{ph}})}{A(r_{\mathrm{ph}})}\,\delta r\right].$ (100)
@@
Substituting $A(r_{\mathrm{ph}})=1/3$, $A^{\prime}(r_{\mathrm{ph}})=2/(9M)$, and Eq. (99) leads to a remarkable cancellation of the $a^{\prime}(r_{\mathrm{ph}})$ terms and yields
$\frac{\delta b_{c}}{b_{c}}=\frac{3}{2}\,a(r_{\mathrm{ph}}).$ (101)
@@
Thus, the leading near-critical modulation of the critical scale depends only on the instantaneous value of $a=\delta A$ at the photon sphere. Since $A=-g_{tt}$, we have $a=-h_{tt}$. Using the RWZ reconstruction in Eq. (68) with $m=0$, $h_{tt}=f,H_{0},Y_{20}(\theta)$; evaluating at the equator $\theta=\pi/2$ and at $r_{\mathrm{ph}}$ with $f=1/3$ gives (equator, $P_{2}$ basis)
$\beta_{1}(\omega)\equiv\frac{\delta b_{c}}{b_{c}}=\frac{1}{4}\,H_{0}(r_{\mathrm{ph}};\omega)$ (102)
@@
In the $Y$-harmonic normalization one has $Y_{20}=\sqrt{\frac{5}{4\pi}}\,P_{2}$, so $\beta_{1}=\frac{1}{4}\,H_{0}^{(P)}=\frac{1}{4}\,\left(\sqrt{\frac{4}{\pi}{5}}\,H_{0}^{(Y)}\right)$; we use the $P_{2}$ convention throughout the main text. The time dependence $\propto\mathrm{Re}{e^{-i\omega t_{o}}\beta_{1}(\omega)}$ then reproduces the $1/\Delta$ term in Eq. (41). Eq. (102) is equivalent to extracting the residue of the $1/\Delta$ pole directly from the Born kernel Appendix B, but makes explicit the quasi-local character of $\beta_{1}$.

### v2

@@ Appendix B Derivation of the master Born kernel
Our goal is to derive the line-of-sight expression that underlies Eqs. (III.3)-(III.7) for the screen-plane deflection produced by a first-order metric perturbation $h_{\mu\nu}$ on the Schwarzschild background of Eq. (II.1). We keep the perturbative split of Eq. (II.3), use the observer tetrad of Eq. (II.9), and map momenta to screen angles via Eq. (II.13). Throughout, we follow a Born, or single-scattering, treatment: the null ray follows the unperturbed geodesic $x_{0}^{\mu}(\lambda;b)$ while its momentum $p^{\mu}$ is corrected at $\mathcal{O}(\varepsilon)$.
@@
$\dot{x}_{0}^{\mu}=\frac{\partial H_{0}}{\partial p_{0\mu}}=g^{(0)\mu\nu}p_{0\nu}\equiv p_{0}^{\mu},\qquad\dot{p}_{0\mu}=-\frac{\partial H_{0}}{\partial x_{0}^{\mu}}=0,$ (B.2)
@@
$\dot{\delta p}_{\mu}=-\frac{\partial}{\partial x^{\mu}}\,\left(\frac{1}{2}\,h^{\alpha\beta}p_{0\alpha}p_{0\beta}\right)=-\frac{1}{2}\,\partial_{\mu}h_{\alpha\beta}\,p_{0}^{\alpha}p_{0}^{\beta},$ (B.3)
which integrates to Eq. (III.2) once evaluated between source $\lambda_{s}$ and observer $\lambda_{o}$:
$\delta p_{\mu}(\lambda_{o};b)=-\frac{1}{2}\int_{\lambda_{s}}^{\lambda_{o}}\,d\lambda\,\partial_{\mu}h_{\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}(\lambda)\,p_{0}^{\beta}(\lambda).$ (B.4)
Eq. (B.4) is the covariant Born response on a fixed background path.
Two comments are useful. First, the force driving $\delta p_{\mu}$ can equivalently be written in terms of the perturbed Christoffel symbols, since
@@
$\delta\Gamma_{\mu\alpha\beta}=\frac{1}{2}\,\left(\nabla_{\alpha}h_{\mu\beta}+\nabla_{\beta}h_{\mu\alpha}-\nabla_{\mu}h_{\alpha\beta}\right),$ (B.5)
so that integrating by parts turns $B.4$ into a $\delta\Gamma$-driven integral plus boundary terms. Second, the contraction $h_{\alpha\beta}p_{0}^{\alpha}p_{0}^{\beta}$ used below is given in (A.8)-(A.11) in terms of the Zerilli master field $\Psi_{Z}$.
@@
$p^{\hat{a}}=e^{\hat{a}}{}_{\mu}p^{\mu}=p_{0}^{\hat{a}}+\varepsilon\,\delta p^{\hat{a}},\qquad\delta p^{\hat{a}}=e^{\hat{a}}{}_{\mu}\,\delta p^{\mu},$ (B.6)
@@
$\alpha_{A}$ $\equiv\frac{A}{r_{o}}=\frac{p^{\widehat{A}}}{p^{\hat{0}}},\quad A\in{X,Y}.$ (B.7)
Varying Eq. (B.7) and keeping terms linear in $\varepsilon$ gives
$\delta\alpha_{A}=\frac{\delta p^{\widehat{A}}}{p_{0}^{\hat{0}}}-\frac{p_{0}^{\widehat{A}}}{(p_{0}^{\hat{0}})^{2}}\,\delta p^{\hat{0}}=\frac{1}{p_{0}^{\hat{0}}}\,\Pi_{A}{}^{\ \hat{b}}\,\delta p_{\hat{b}}$, (B.8)
@@
$\Pi_{A}{}^{\ \hat{b}}=\delta_{A}{}^{\ \hat{b}}-\alpha_{0,A}\,\delta^{\hat{b}}{}_{\hat{0}}\,,\qquad\alpha_{0,A}=\frac{p_{0}^{\widehat{A}}}{p_{0}^{\hat{0}}}$, (B.9)
and hatted indices are raised/lowered with $\eta_{\hat{a}\hat{b}}$. Using $\delta p_{\hat{b}}=e_{\hat{b}}{}^{\mu}\,\delta p_{\mu}$ and substituting Eq. (B.4) then yields
@@
$\times\partial_{\mu}h_{\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}(\lambda)p_{0}^{\beta}(\lambda),$ (B.10)
which is Eq. (III.4) with the explicit kernel
@@
$\times\partial_{\mu}h_{\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}p_{0}^{\beta}.$ (B.11)
Eq. (B.11) is identical to Eq. (III.5) once we identify the tetrad time leg $e^{\nu}{}_{\hat{0}}$ and note that $\Pi_{A}{}^{\ \mu}=\Pi_{A}{}^{\ \hat{b}}e_{\hat{b}}{}^{\ \mu}$.
Equivalent kernel forms and boundary terms: Using the identity Eq. (B.5) and integrating by parts, Eq. (B.10) becomes
$\delta\alpha_{A}=\int_{\lambda_{s}}^{\lambda_{o}}\,d\lambda\,\mathcal{K}^{(\Gamma)}_{A}+\mathcal{B}_{A}\Big|_{\lambda_{s}}^{\lambda_{o}},$ (B.12)
@@
$\mathcal{K}^{(\Gamma)}_{A}=-\frac{1}{p_{0}^{\hat{0}}(\lambda_{o})}\,\Pi_{A}{}^{\ \hat{b}}\,e_{\hat{b}}{}^{\ \mu}(x_{o})\,\delta\Gamma_{\mu\alpha\beta}\left(x_{0}(\lambda)\right)\,p_{0}^{\alpha}p_{0}^{\beta},$ (B.13)
$\mathcal{B}_{A}=\frac{1}{2\,p_{0}^{\hat{0}}(\lambda_{o})}\,\Pi_{A}{}^{\ \hat{b}}\,e_{\hat{b}}{}^{\ \mu}(x_{o})\,h_{\mu\nu}\left(x_{0}(\lambda)\right)\,p_{0}^{\nu}.$ (B.14)
@@
The boundary term $\mathcal{B}_{A}$ vanishes under our asymptotic conditions: at the observer, the screen projector eliminates any $\hat{0}$ contamination and $h_{\mu\nu}(x_{o})=\mathcal{O}(r_{o}^{-1})$ [Eq. (A.7)]; at the source, we assume compact support or sufficiently rapid falloff. Hence, all kernel representatives obtained by redistributing derivatives are equivalent at $\mathcal{O}(\varepsilon)$.
@@
A further useful form follows by changing variables from affine parameter $\lambda$ to the tortoise coordinate $r_{*}$ or to the delay $\tau$ of Eq. (III.7). Using Eq. (A.11),
@@
$\delta\alpha_{A}=\int_{\tau=0}^{\infty}\,d\tau\,\widetilde{\mathcal{K}}_{A}\left[r(\tau),\theta(\tau)\right]\,e^{i\omega\tau}\times e^{-i\omega t_{o}}+\text{c.c.},$ (B.15)
@@
which is the factorized form quoted in Eq. (III.7) once the QNM time dependence in Eq. (III.6) is inserted.
@@
Examining the gauge behavior at $\mathcal{O}(\varepsilon)$ under a linearized diffeomorphism generated by $\xi^{\mu}$, Eq. (II.15) gives $\delta h_{\alpha\beta}=-\nabla_{\alpha}\xi_{\beta}-\nabla_{\beta}\xi_{\alpha}$. Substituting this into Eq. (B.10) and integrating by parts yields
@@
$\delta_{\xi}\alpha_{A}=\frac{1}{p_{0}^{\hat{0}}(\lambda_{o})}\,\Pi_{A}{}^{\ \hat{b}}\,e_{\hat{b}}{}^{\ \mu}(x_{o})\,\xi_{\mu}\Big|_{\lambda_{s}}^{\lambda_{o}}+\mathcal{O}\,\left(\frac{\varepsilon}{r_{o}}\right),$ (B.16)
@@
where we used the background geodesic equation and $\dot{p}_{0\mu}=0$. With our asymptotic conditions $\xi^{\mu}=\mathcal{O}(1/r)$, Eq. (B.16) reproduces the gauge-independence statement in Eq (II.16): $\delta_{\xi}\alpha_{A}=0+\mathcal{O}(\varepsilon/r_{o})$. Thus, the kernel Eq. (B.11) produces gauge-safe screen observables at first order, provided the observer tetrad is fixed by Eq. (II.10).
@@
In the far-zone reduction implies the $1/b$ law. Insert the large-$r$ asymptotics $A.12$ into $B.11$, parametrize the background ray by a straight line with longitudinal coordinate $z$ as in Eq. (III.9), and note that $p_{0}^{\hat{0}}(\lambda_{o})=E$ to leading order. One finds
@@
$\delta\alpha_{A}(b,t_{o})=\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\frac{c_{A}(\omega)}{b}\,\mathcal{F}_{A}(\omega b)\right\}+\mathcal{O}\,\left(\frac{M}{b^{2}}\right),$ (B.17)
@@
with $c_{A}$ and $\mathcal{F}_{A}$ defined in Eqs. (III.10)-(III.11). This reproduces the weak-field Born expansion used in Sec. III.2 and provides the link between the kernel normalization and the centroid wobble of Eq. (III.15).
@@
Near-photon-sphere reduction implies the logarithmic law. In the near-critical regime $b\to b_{c}^{+}$, reparametrize the background ray by the azimuth $\varphi$ accumulated during the whirl [Eq. (IV.4)]. Use the near-photon-sphere reduction $A.13$ inside $B.11$ and integrate over the long dwell time $\Delta\varphi\sim-\ln\Delta$, with $\Delta=b/b_{c}-1$. The integral yields
@@
$\delta\alpha_{A}(b,t_{o})=\mathrm{Re}\left\{e^{-i\omega t_{o}}\,\left[\mathcal{C}_{A}(\omega)\,\ln\Delta+\mathcal{D}_{A}(\omega)\right]\right\}+\ldots,$ (B.18)
@@
which is Eq. (IV.5). When combined with the background divergence in Eq. (IV.2) and reorganized as in Eqs. (IV.6)-(IV.9), Eq. (B.18) fixes the time-dependent SDL coefficients $a_{1},b_{1},\beta_{1}$ from the same kernel.
@@
Starting from Hamilton’s equations, we derived the Born response (B.4) for the photon momentum, projected it to the observer’s screen to obtain the master kernel (B.11), and exhibited equivalent $\delta\Gamma$ and delay-space forms, Eqs. (B.14)-(B.15). The kernel is gauge-safe at $\mathcal{O}(\varepsilon)$ by Eq. (B.16). Its far-zone and near-photon-sphere reductions, Eqs. (B.17) and (B.18), reproduce the weak-field $1/b$ behavior Sec. III.2 and the logarithmic SDL structure Sec. IV, respectively. In this sense, all imaging diagnostics discussed in Sec. V are different asymptotic faces of a single line-of-sight kernel driven by the Zerilli master field.
@@ Appendix C Near-critical expansion and SDL coefficients
In this appendix, we provide a self-contained derivation of the near-critical expansion underlying Eqs. (IV.2) and (IV.6)-(IV.9), and we express the time-dependent strong-deflection-limit $SDL$ response coefficients $a_{1}(\omega),b_{1}(\omega),\beta_{1}(\omega)$} in terms of quantities localized at the photon sphere and of moments of the Born kernel introduced in Appendix B. We keep all conventions of Secs. II-IV, in particular the Schwarzschild background (Eq. (II.1)), the perturbative split (Eq. (II.3)), and the master Born representation (Eqs. (III.3)-(III.7)).
@@
$\mathcal{F}(r)\equiv r\,A^{\prime}(r)-2A(r)=0,\qquad r=r_{\mathrm{ph}},$ (C.2)
and the critical impact parameter is $b_{c}=r_{\mathrm{ph}}/\sqrt{A(r_{\mathrm{ph}})}$. For Schwarzschild, A=f=12M/rA=f=1-2M/r gives $r_{\mathrm{ph}}=3M$ and $b_{c}=3\sqrt{3}M$ as stated in Eq. (IV.1). Expanding the background bending integral in Eq. (III.1) around $b\to b_{c}^{+}$ yields Eq. (IV.2),
$\alpha_{0}(b)=-\bar{a}\,\ln\,\left(\frac{b}{b_{c}}-1\right)+\bar{b}+\mathcal{O}\,\left((b-b_{c})\ln|b-b_{c}|\right),$ (C.3)
where $\bar{b}$ is a constant depending on the chosen normalization, and $\bar{a}=1$ for the Schwarzschild case.
Linear response of the critical scale $b_{c}(t_{o})$: We now compute the $\mathcal{O}(\varepsilon)$ modulation of $b_{c}$ induced by the even-parity $(\ell,m)=(2,0)$ perturbation described in Sec. II.1. We adopt an instantaneous snapshot characterization at the detection time $t_{o}$: the time dependence appears as a harmonic factor $e^{-i\omega t_{o}}$ multiplying complex amplitudes evaluated at $r_{\mathrm{ph}}$. This adiabatic reading is equivalent, to first order, to the near-photon-sphere reduction of the Born kernel and provides a compact route to $\beta_{1}(\omega)$.
Let $A(r)\to A(r)+\varepsilon\,a(r,t)$, where $A=f$ for the background and $a$ is the even-parity perturbation of $g_{tt}$ (we keep $B$ unmodified since, to first order in $\varepsilon$, $b_{c}$ is controlled by $A$ alone through Eq. (C.1)). The photon-sphere condition (C.2) becomes $\mathcal{F}(r_{\mathrm{ph}}+\varepsilon\,\delta r)=0$ with
$\delta r\,\mathcal{F}^{\prime}(r_{\mathrm{ph}})+\left[r\,a^{\prime}(r)-2a(r)\right]_{r=r_{\mathrm{ph}}}=0,$ (C.4)
$\mathcal{F}^{\prime}(r_{\mathrm{ph}})=r_{\mathrm{ph}}A^{\prime\prime}(r_{\mathrm{ph}})-A^{\prime}(r_{\mathrm{ph}}).$ (C.5)
Solving Eq. (C.5) gives
@@
$=-\frac{3M}{2}\,\left[2\,a(r_{\mathrm{ph}})-r_{\mathrm{ph}}\,a^{\prime}(r_{\mathrm{ph}})\right].$ (C.6)
@@
$\frac{\delta b_{c}}{b_{c}}=\frac{\delta r}{r_{\mathrm{ph}}}-\frac{1}{2}\left[\frac{a(r_{\mathrm{ph}})}{A(r_{\mathrm{ph}})}+\frac{A^{\prime}(r_{\mathrm{ph}})}{A(r_{\mathrm{ph}})}\,\delta r\right].$ (C.7)
@@
Substituting $A(r_{\mathrm{ph}})=1/3$, $A^{\prime}(r_{\mathrm{ph}})=2/(9M)$, and Eq. (C.6) leads to a remarkable cancellation of the $a^{\prime}(r_{\mathrm{ph}})$ terms and yields
$\frac{\delta b_{c}}{b_{c}}=\frac{3}{2}\,a(r_{\mathrm{ph}}).$ (C.8)
@@
Thus, the leading near-critical modulation of the critical scale depends only on the instantaneous value of $a=\delta A$ at the photon sphere. Since $A=-g_{tt}$, we have $a=-h_{tt}$. Using the RWZ reconstruction in Eq. (A.4) with $m=0$, $h_{tt}=f,H_{0},Y_{20}(\theta)$; evaluating at the equator $\theta=\pi/2$ and at $r_{\mathrm{ph}}$ with $f=1/3$ gives (equator, $P_{2}$ basis)
$\beta_{1}(\omega)\equiv\frac{\delta b_{c}}{b_{c}}=\frac{1}{4}\,H_{0}(r_{\mathrm{ph}};\omega)$ (C.9)
@@
In the $Y$-harmonic normalization one has $Y_{20}=\sqrt{\frac{5}{4\pi}}\,P_{2}$, so $\beta_{1}=\frac{1}{4}\,H_{0}^{(P)}=\frac{1}{4}\,\left(\sqrt{\frac{4}{\pi}{5}}\,H_{0}^{(Y)}\right)$; we use the $P_{2}$ convention throughout the main text. The time dependence $\propto\mathrm{Re}{e^{-i\omega t_{o}}\beta_{1}(\omega)}$ then reproduces the $1/\Delta$ term in Eq. (IV.9). Eq. (C.9) is equivalent to extracting the residue of the $1/\Delta$ pole directly from the Born kernel Appendix B, but makes explicit the quasi-local character of $\beta_{1}$.